OpenStudy (anonymous):

solve 125x^3 + 343 = 0

OpenStudy (anonymous):

x = -1.4 subtract 343 on both sides and diide by 125 and take the cube root

OpenStudy (anonymous):

it isnt 1.4 though

OpenStudy (mr.math):

\[125x^3 + 343 = 0 \implies x^3=-\frac{343}{125} \implies x=-\frac{7}{5}\]

OpenStudy (mr.math):

This is the only real solution.

OpenStudy (anonymous):

yea i dont see how it can be anything else

OpenStudy (anonymous):

mr. math, did you divide 125 by 343

OpenStudy (mr.math):

@kanade: Actually it can. This equation has exactly three distinct roots over complex numbers.

OpenStudy (mr.math):

@kamc23: I subtracted 343 from both sides and then divided by 125.

OpenStudy (anonymous):

but when i divide 343 by 125 i get 2.744

OpenStudy (anonymous):

if you take the cubed root of 2.744 is it 3 distinct roots due to the triangle created in the real and imaginary plane?

OpenStudy (anonymous):

..?

OpenStudy (mr.math):

Yes @kanade

OpenStudy (anonymous):

i dont know how to take the cubed root of 2.744

OpenStudy (mr.math):

Put 2.744^(1/3) in your calculator. :D

OpenStudy (anonymous):

you could take the cube root of 343 and 125 individually and then divide to end up with 7/5 which you can get 1.4 then or you'd prob need to use a calculator

OpenStudy (anonymous):

ughh i still dont get it